The sum of three consecutive odd integers is 189. What are the integers?
Answer by Alan3354(69443)

Do it like this one:
===================
The sum of three consecutive odd integers is 165. What are the three integers?
--------------
3 + 5 + 7 = 15 That's not it.
17 + 19 + 21 = 63 Not it.
31 + 33 + 35 = 99 Not it.
-------------
Notice that the sum is 3 times the middle integer?